OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 20 miles per gallon of gas and the second has a fuel efficiency of 25 miles per gallon of gas. During one particular week, the two cars went a combined total of 1200 miles, for a total gas consumption of 55 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

let x=gas consumed by 1st car 40-x =gas consumed by 2nd car distance traveled=fuel efficiency*gas consumed .. 20x+25(40-x)=925 20x+1000-25x=925 5x=75 x=15 40-x=25 gas consumed by 1st car=15 gal gas consumed by 2nd car=25 gal

OpenStudy (anonymous):

gallons were consumed by each of the two cars that week are 40 and 15 Solution:Set the miles of the first car is X, the gallons consumed is Y then (55-Y)x 20 =1300- X Y x 25 = X so Y = 40, X = 1000